### v2

# Comparison of Lumerical FDTD and Tidy3D for three-dimensional FDTD simulations of passive silicon photonic components

Comparison of Lumerical FDTD and Tidy3D for three-dimensional FDTD simulations of passive silicon photonic components
††journal: opticajournal††articletype: Research ArticleWe benchmark Lumerical FDTD and Tidy3D for 3D simulations of passive silicon photonic components on the silicon-on-insulator (SOI) platform. Six devices – including an MMI, directional coupler, waveguide crossing, mode converter, polarization splitter rotator, and ring resonator – are simulated under matched conditions using geometries from the generic process design kit (PDK) from GDSFactory. Our study emphasize on comparing simulation accuracy across both solvers, alongside an analysis of runtime and broadband behavior over varying resolutions and bandwidths. Results show that both solvers are reliable with minimal discrepancies.
1 Introduction
The finite-difference time-domain (FDTD) method is a widely used full-wave solution for modeling eletromagnetic fields in complex media and geometries [1, 2, 3]. It discretizes the simulation domain into a mesh and solves Maxwell’s equations at each point over time. As a time-domain method, FDTD naturally supports broadband analysis and allows accurate modeling of arbitrarily shaped, dispersive, and subwavelength structures. These capabilities make it especially valuable for simulating integrated photonic devices with high index contrast and complex geometries.
Due to its versatility, FDTD is employed in the simulation of photonic components including waveguides, resonators, and grating couplers [4, 5, 6, 7, 8, 9]. However, many devices require full three-dimensional (3D) simulations to capture asymmetric vertical structure and polarization effects, which significantly increase computational demands. Simulation time and memory usage scale rapidly with discretization. As a result, a careful balance between accuracy, resolution, and runtime is essential.
Among these tools, Lumerical FDTD (Ansys) is a widely used commercial solver known for its robust performance and various built-in features with a rich desktop graphical user interface (GUI) [10]. Meanwhile, Tidy3D (Flexcompute) is an emerging cloud-native FDTD solver that offers graphics processing unit (GPU) acceleration, a Python interface for streamlined workflow, and a web-based GUI [11]. They differ significantly in computational performance, runtime, and user experience. However, to our knowledge, there have not been a study that systematically compares these electromagnetic solvers under matched simulation conditions.
In this study, we benchmark Lumerical FDTD and Tidy3D for 3D FDTD simulations of a series of passive silicon photonic components on the silicon-on-insulator (SOI) platform in terms of simulation accuracy, broadband simulation reliability, runtime, and usability. We consider a representative set of devices including a 22 multimode interferometer (MMI), a directional coupler, a waveguide crossing, a mode converter, a polarization rotator splitter (PRS), and a ring resonator. All geometries are picked from the generic process design kit (PDK) of GDSFactory, an open source electronics design automation (EDA) tool for integrated circuits. [12]. Each device is simulated using the same material models, boundary conditions, and source configurations across both solvers. The simulations are centered at 1550 nm.
The remainder of this manuscript is structured as follows. Section 2 describes the simulation setup and configurations. Section 3 presents the simulation results and comparative analysis. Section 4 discusses the implications of our findings in practical photonic design workflows. Finally, Section 5 concludes the paper.
2 Simulation configurations
Simulation of the same set of devices are conducted using Lumerical FDTD and Tidy3D under matched settings for comparison. The Lumerical FDTD solver version is 2025 R1 (8.33.3999), running locally using CPU on a workstation with AMD 3960X 24-core processor, 64 GB RAM, and Windows 10 OS. Tidy3D simulations are executed on Flexcompute’s GPU-accelerated cloud cluster using version 2.7.7 [13]. The computation resources are dynamically allocated based on simulation size and hardware availability at the time. After inquiry with Flexcompute, we were informed that one representative simulation - mode converter at the resolution of 25 cells per wavelength - was allocated to eight NVIDIA A100 GPUs. It needs to be noted that direct wall-clock comparisons reflect the best-effort use of each platform rather than identical compute resources. For both solvers, the geometry and simulation domain setup, boundary settings, and mesh configuration are all handled through their Python application programming interfaces (APIs). The Python codes are included in [14]. The auto-shutoff threshold is set to 110-5 for early termination. The simulation domain was defined to enclose the component with a distance of 2 m from its top and bottom edges and 1 m from edges on the device plane. Each device is extended with 10 m long straight waveguides at every port before importing into the solvers. This preprocessing step is to ensure extrusion of input and output waveguides through the boundaries in the propagation direction and therefore a reliable result. Perfectly matched layer (PML) boundary condition is applied to all six sides of the 3D domain. No symmetry planes are used to maintain generality. If a diverging simulation is encountered, we change the boundary condition accordingly. For instance, we switch to “absorber” in Tidy3D, which applies adiabatic absorbers containing multiple layers with gradually increasing conductivity to the boundaries. It is suitable for dispersive materials intersecting with simulation boundaries [15].
We selected five passive silicon photonic components to span an range of mode coupling mechanisms and structure complexity. All components are designed for a wavelength around 1550 nm and sourced directly from the generic PDK of GDSFactory. To study the interactions of fundamental modes, we include a directional coupler and a waveguide crossing. A 22 MMI is simulated to evaluate multimode propagation. Additionally, a mode converter is modeled for higher-order mode excitation. We also select a polarization splitter rotator (PSR) to study higher-order mode propagation and polarization handling. Finally, we simulate a ring resonator for more complex resonance behavior. We generate GDSII files of the components and import them to the solvers through Python APIs.
The layer stack consists of a standard SOI platform with 220 nm silicon (Si) in top and bottom cladding of silicon oxide (SiO2). One exception is the PSR, where the top cladding is silicon nitride (Si3N4) and the bottom cladding is SiO2, as demonstrated by Dai et al. [16] The dispersion of Si and SiO2 are obtained from Handbook of Optical Constants of Solids by E. Palik et al. [17] The refractive index of Si3N4 is estimated as 2.0 around the center wavelength of 1550 nm. On one hand, Lumerical FDTD models dispersive materials using multi-coefficient material (MCM) model, which can describe material dispersion with higher accuracy and lower computational need [18]. It fits the frequency-dependent permittivity of Si with two coefficients, resulting in a root mean square (RMS) error of 2.1510-5, without disclosing the value of fitting coefficients. Also, the permittivity of SiO2 is fitted as a constant of 2.09 in the simulation bandwidth with a RMS error of 2.4310-4. On the other hand, Tidy3D fits the dispersion data using the pole-residue model [19]. This model expresses the frequency-dependent permittivity as a sum of resonant material poles:
where represents the frequency, the relative permittivity at infinite frequency, and the complex pole-residue pairs. Table 1 lists the fitted relative permittivity and pole-residue pairs for Si and SiO2, with RMS of 1.5010-2 and 8.6310-6, respectively.
Material | Coefficients | Values |
Si | 10.67 | |
SiO2 | 1.28 | |
To evaluate the resolution-dependent performance, we applied different discretization sizes in both solvers. The spatial resolution is determined by the number of cells per wavelength, which is the wavelength in the highest-index medium in the simulation domain. We consider a series of different resolutions in the comparison. No mesh override is manually applied. In Lumerical FDTD, we choose the custom non-uniform mesh and update the minimum mesh cells per wavelength with the selected spatial resolution. The meshing algorithm applies a smaller mesh in high index materials. Similarly, Tidy3D provides an automatic non-uniform meshing option where the resolution is specified by the minimum steps per wavelength in each material. Note that the automatic grid conforms to the structures in Tidy3D, which helps avoiding staircasing of permittivity at the edges. Lumerical FDTD offers conformal mesh technology (CMT) that can account for subcell features near the edges [20].
Each device is excited using a mode source with bandwidth of 20 nm or 50 nm centered at 1550 nm. The input is fundamental TE mode, except in the PRS where fundamental TM mode is considered to evaluate polarization evolution. In Lumerical FDTD, we insert port objects at each input and output port of the devices as mode source or field and mode expansion monitor. Accurate modeling of the source mode across the wavelength range is ensured by enabling frequency dependent profiles in the source port. In Tidy3D, a mode source is placed at the input port and mode monitors are placed at output ports. Additionally, field monitors are inserted in the device plane to visualize electric field distribution. The source spectra and positions of the source and monitors are kept the same in both solvers.
3 Simulation results and comparisons
3.1 Directional coupler
The directional coupler is sourced from the coupler component in GDSFactory generic PDK. As sketched in Fig. 1, it consists of two 500 nm wide, 220 nm high Si waveguides placed in close proximity within the coupling region that is 20 m long. The gap width is 236 nm. The component is cladded with SiO2 on the top and the bottom. Fig. 2 shows the device imported into Lumerical FDTD and Tidy3D including a 3D simulation domain, monitors on all waveguides, a 2D field monitor on the device plane, and a mode source on the lower left waveguide. The mode source inject fundamental TE mode at 1550 nm into the waveguide. Fig. 3 presents the normalized electric field intensity on the device plane computed by both solvers at a high resolution of 25 cells per wavelength. As expected, a fraction of the power is coupled to the top waveguide and propagates to the cross port. The two solvers show good agreement in the general distribution of the electric field and the intensity to each output port.
The component is simulated at a series of resolutions ranging from 6 to 25 cells per wavelength. We notice that Tidy3D needs the boundary conditions to be set as “absorber” on all boundaries to avoid diverging simulation when the spatial resolution is 20 and 25. Table 2 lists the runtime in both solvers and the peak memory usage in Lumerical FDTD, with a source bandwidth of 20 nm centered at 1550 nm. The peak memory usage in Lumerical FDTD scales rapidly with increasing resolution due to finer discretization. With the same size of the simulation domain (4274 m3), the runtime of Tidy3D is much shorter than that of Lumerical FDTD. For example, Tidy3D completes the simulation with a resolution of 25 in roughly 10 minutes, while Lumerical FDTD needs almost 81 minutes. For this resolution, the total number of grid points and the elapsed simulation time and iterations are listed in Table 3. Tidy3D shows slightly finer discretization and runs for almost three times longer simulation time.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 17 | 148 | 0.35 |
10 | 35 | 336 | 0.92 |
15 | 66 | 1293 | 2.42 |
20 | 136 | 2606 | 5.01 |
25 | 617 | 4856 | 9.07 |
Tidy3D | Lumerical FDTD | |
Total number of grids | 1.77108 | 9.68107 |
Number of grids in x | 2434 | 2391 |
Number of grids in y | 395 | 332 |
Number of grids in z | 184 | 122 |
Elapsed simulation time (s) | 2.0610-12 | 7.0510-13 |
Elapsed iterations | 61812 | 21303 |
Then we examine the simulation result of the two solvers at different spatial resolutions. Fig. 4(a) presents the transmission to the fundamental TE mode in the cross port at 1550 nm as a function of resolution. These results are computed using a source bandwidth of 20 nm. Both solvers overestimate the transmission at lower resolutions (6 and 10 cells per wavelength) and converge to a more stable value at the resolution of 15 cells per wavelength. Results of Tidy3D shows a more rapid decrease from 69.7 at 6 cells per wavelength to 44.8 at 15 cells per wavelength. Meanwhile, Lumerical FDTD decreases from 49.3 at 6 cells per wavelength to 41.1 at 15 cells per wavelength. Both solvers exhibit some fluctuations at higher resolutions beyond 15 cells per wavelength, with that in Tidy3D below 4.4 and Lumerical FDTD below 3.6. The difference between the two solvers for the higher resolutions is up to 4.6. These results indicate that low resolution simulations from Tidy3D is less reliable, while both solvers converge well at a moderate resolution of 15 cells per wavelength with the exact values differing slightly.
The simulations are then conducted with different source bandwidths to examine the solvers’ ability of broadband simulations. Fig. 4(b) presents the transmission spectra from the two solvers with source bandwidths of 20 nm and 50 nm at a resolution of 15 cells per wavelength. Both solvers capture the increasing transmission with wavelength. The results from Tidy3D demonstrate good internal consistency with a deviation around 0.1 between the two bandwidths. In contrast, Lumerical FDTD gives a noticeably lower transmission at the bandwidth of 20 nm with an offset of 3 from its 50 nm result. The spectrum with 50 nm bandwidth is closer to the Tidy3D spectra and slightly lower than them by 0.7. It suggests that Tidy3D is consistent across bandwidth settings, while extra attention should be taken when interpreting results with different bandwidths in Lumerical FDTD.
Fig. 4(c) shows the excess loss computed for different spatial resolutions, which quantifies the intensity transmitted to the fundamental TE modes in neither of the two output ports. Similar to the trends shown in Fig. 4(a), both solvers converge to a more stable value around 15 cells per wavelength, and Tidy3D shows a sharper increase when the resolution is increased from 6 to 10 cells per wavelength. It is demonstrated in Fig. 4(d) that when the source bandwidth is increased from 20 nm to 50 nm, both solvers show good internal agreement and a very small cross-solver discrepancy of around 0.01 dB.
3.2 Waveguide crossing
The waveguide crossing is identical to the crossing component in GDSFactory generic PDK. As sketched in Fig. 5, it has a 1.2 m wide Si square at the center, of which each side is joined by a 3.4 m long taper that connects to a 500 nm wide waveguide. The square, tapers, and waveguides are 220 nm thick. In the slab layer that is 150 nm thick, there are two identical ellipses sharing the center of the crossing with their major axis orthogonal to each other. The major axis is 6 m long and the minor axis is 2.2 m. This component has SiO2 as top and bottom cladding. Fig. 6 demonstrates the waveguide crossing as seen in Lumerical FDTD and Tidy3D. A fundamental TE mode is injected in the left arm at 1550 nm. Fig. 7 presents the electric field computed by the two solvers with a resolution of 25 cells per wavelength. Both solvers show low-loss transmission to the through port and agree well in field patterns in the crossing section.
We examine simulations of this waveguide crossing in a series of resolutions from 6 to 25 cells per wavelength with a source bandwidth of 20 nm. The runtime and peak memory usage are listed in Table 4. The memory usage is lower than that for directional coupler simulations due to a smaller simulation domain. As expected, Tidy3D is much faster than Lumerical FDTD run on our workstation. With a simulation domain of 10114 m3 and a resolution of 25 cells per wavelength, Tidy3D completes the simulation in 1 minute and Lumerical completes it in 21 minutes. With this resolution, the total number of grid points and the elapsed simulation time and iterations are listed in Table 5. Both solvers show similar discretization, yet Tidy3D elapsed almost two times longer simulation time before the simulation reaches the auto-shutoff value.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 11 | 31 | 0.23 |
10 | 18 | 85 | 0.51 |
15 | 28 | 234 | 1.26 |
20 | 84 | 591 | 2.57 |
25 | 61 | 1268 | 4.55 |
Tidy3D | Lumerical FDTD | |
Total number of grids | 4.83107 | 4.55107 |
Number of grids in x | 587 | 584 |
Number of grids in y | 643 | 638 |
Number of grids in z | 128 | 122 |
Elapsed simulation time (s) | 6.8310-13 | 3.5910-13 |
Elapsed iterations | 20659 | 10843 |
We then compare the transmission to the fundamental TE mode in the through port at different spatial resolutions from 6 to 25 cells per wavelength from the two solvers, as presented in Fig. 8(a). It shows the transmission at 1550 nm computed with a source bandwidth of 20 nm. The results from Lumerical FDTD is stable across the resolution range, increasing slightly from 95.4 at 6 and 10 cells per wavelength to 95.7 at resolutions not smaller than 15 cells per wavelength. In contrast, Tidy3D increases more sharply from 93.9 at 6 cells per wavelength to a peak of 95.9 at 15 cells per wavelength, followed by a slight decline at high resolutions to meet the value of 95.7 at 25 cells per wavelength. These results indicate that both solvers converge to within 0.3 agreement by 15 cells per wavelength, though they exhibit different convergence behaviors. These results indicate that the two solvers converge to within 0.3 agreement by 10 cells per wavelength in Tidy3D and 15 cells per wavelength in Lumerical FDTD. The lower resolution simulations from Tidy3D shows more deviation. Similar trends can be seen in the comparison of excess loss in Fig. 8(c), which quantifies the amount of power that is not coupled to the fundamental TE mode of any of the three output ports.
Fig. 8(b) shows the transmission spectra to the fundamental TE mode in the through port with different source bandwidths in the two solvers at a resolution of 15 cells per wavelength. Both solvers exhibit good internal consistency, with internal differences below 0.11 and 0.04 for Lumerical FDTD and Tidy3D, respectively. The results of Lumerical FDTD is slightly lower than that of Tidy3D by an offset of 0.13. Tidy3D demonstrates a peak of 96.15 at the wavelength of 1.565 m, while Lumerical FDTD reaches a plateu of 96.00 around that wavelength and increases again. Overall, the two solvers agree well across different source bandwidths at the resolution of 15 cells per wavelength. Again, similar trends can be seen from the comparison of excess loss from the two solvers in Fig. 8(d).
3.3 22 MMI
The MMI has the default layout provided by the mmi2x2withsbend component in GDSFactory generic PDK. This layout is proposed by Guan et al. in [21]. The component is defined in 220 nm thick Si cladded in SiO2. The input and output waveguides are 500 nm wide and tapered to 700 nm wide before the multimode region in a length of 1 m, as sketched in Fig. 9. Fig. 10 demonstrates the MMI imported to Lumerical FDTD and Tidy3D. A fundamental TE mode at 1550 nm is injected at the lower left arm. Examples of electric field intensity from the two solvers are presented in Fig. 11. They are computed at a resolution of 25 cells per wavelength. The two solvers agree well in the multimode interference pattern and intensity distribution in the two output waveguides.
The MMI is examined in both solvers with the same set of resolutions as in the previous sections with a source bandwidth of 20 nm centered at 1550 nm. Table 6 lists the runtime of the both solvers and the peak memory usage in Lumerical FDTD. As expected, Tidy3D is faster than Lumerical FDTD, completing the simulation with 25 cells per wavelength in 2 minutes 30 seconds while Lumerical needs 80 minutes. The simulation domain size is 347.54 m3. Table 7 lists the total number of grid points and the elapsed simulation time and iterations with the resolution of 25. Both solvers show similar discretization while Tidy3D elapsed a 30 longer simulation time.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 11 | 98 | 0.32 |
10 | 34 | 259 | 0.82 |
15 | 45 | 905 | 2.15 |
20 | 57 | 2182 | 4.48 |
25 | 150 | 4813 | 8.10 |
Tidy3D | Lumerical FDTD | |
Total number of grids | 9.06107 | 8.59107 |
Number of grids in x | 1929 | 1939 |
Number of grids in y | 367 | 363 |
Number of grids in z | 128 | 122 |
Elapsed simulation time (s) | 1.1410-12 | 7.9210-13 |
Elapsed iterations | 34262 | 23928 |
Fig. 12(a) compares the transmission to the fundamental TE mode at the cross port computed by the two solvers as a function of spatial resolution. The results are computed using a source bandwidth of 20 nm, and only the transmission at 1550 nm is displayed. The overall difference between the two solvers is less than 2. In Lumerical FDTD, the transmission increases from 37.6 at 6 cells per wavelength to a more stable value around 48 at 15 cells per wavelength and beyond. Similarly, Tidy3D increases from 35.8 at 6 cells per wavelength to 47.9 at 15 cells per wavelength. A moderate resolution of 15 cells per wavelength is sufficient in both solvers to provide reliable results. The good agreement and moderate convergence resolution can also be found in the quantification of excess loss as presented in Fig. 12(c).
Then we examine the performance of the two solvers with different source bandwidths. Fig. 12(b) shows the transmission spectra to the fundamental TE mode at the cross port with bandwidths of 20 nm and 50 nm at a resolution of 15 cells per wavelength. Both solvers exhibit excellent internal consistency across bandwidth settings and a similar decreasing transmission with increasing wavelength. The results from Lumerical FDTD is higher than that of Tidy3D by less than 0.72. The negligible discrepancy also applies to the excess loss presented in Fig. 12(d).
3.4 Mode converter
The mode converter is sourced from the modeconverter component in GDSFactory generic PDK, as proposed in [22]. The function is to convert the fundamental TE mode in a single-mode waveguide into the first order TE mode in a wider waveguide. It consists of a 500 nm wide waveguide and a 1 m wide waveguide in close proximity within a 20 m long coupling section. The gap width is 150 nm. The 1 m wide waveguide is tapered to 1.2 m wide before and after the coupling section by 25 m long tapers. All waveguides are made of 220 nm thick Si and cladded in SiO2. Fig. 13 illustrates the mode converter and Fig. 14 shows this component imported into Lumerical FDTD and Tidy3D. Fig. 15 includes examples of the electric field intensity computed by Lumerical FDTD and Tidy3D at a resolution of 25 cells per wavelength. Both solvers capture that a small fraction of the electric field is coupled from the narrow input waveguide to the wider output waveguide as a higher order mode. It needs to be noted that Tidy3D requires the boundary condition to be “absorber” on all boundaries to avoid simulation diverging for the simulations of this component.
The simulation is run at a series of spatial resolutions ranging from 6 to 25 cells per wavelength with a source bandwidth of 20 nm, as in the previous sections. The runtime in both solvers and the peak memory usage is listed in Table 8. Similar to the simulation of previous components, Tidy3D completes the job in much shorter time. For instance, Tidy3D finishes the simulation with the highest resolution (25 cells per wavelength) in 33 minutes while Lumerical FDTD needs 25 hours with a peak memory usage of 23.63 GB. We notice that the runtime of Tidy3D does not increase monotonically with spatial resolution, as the computational speed depends on the server resource usage at the time. The memory usage is more demanding than that in simulations of previous components due to a larger simulation domain (7210.54 m3). With the resolution of 25, the total number of grid points and the elapsed simulation time and iterations are listed in Table 9. Tidy3D shows finer discretization and elapses slightly longer simulation time.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 63 | 760 | 0.71 |
10 | 517 | 3599 | 2.15 |
15 | 415 | 14417 | 6.00 |
20 | 124 | 40710 | 12.90 |
25 | 1997 | 91484 | 23.63 |
Tidy3D | Lumerical FDTD | |
Total number of grids | 4.37108 | 2.58108 |
Number of grids in x | 4115 | 4084 |
Number of grids in y | 577 | 518 |
Number of grids in z | 184 | 122 |
Elapsed simulation time (s) | 5.5310-12 | 5.1810-12 |
Elapsed iterations | 165849 | 156383 |
We compare the transmission to the first order TE mode at the cross port across the spatial resolutions from the two solvers, as presented in Fig. 16(a). It considers the transmission at 1550 nm computed with a source bandwidth of 20 nm. In Lumerical FDTD, the transmission drops sharply from an unphysical value of 96.7 at 6 cells per wavelength to more realistic values around 50 for higher resolutions above 10 cells per wavelength. In contrast, the transmission from Tidy3D increases gradually from 35.7 at 6 cells per wavelength to 49.2 at 15 cells per wavelength. The two solvers agree well between 15 and 20 cells per wavelength, with discrepancies below 5. These results suggest that Lumerical FDTD converges to approximately 50 at a slightly lower resolution of 10 cells per wavelength, while Tidy3D converges to a similar value at 15 cells per wavelength. Similar trends can be seen from Fig. 16(c) showing the mode crosstalk, which quantifies the transmission from the source mode to the fundamental TE mode in the cross port.
The bandwidth-dependent performance is evaluated by comparing transmission spectra to the first order TE mode at the cross port with 20 nm and 50 nm source bandwidths in both solvers at a resolution of 15 cells per wavelength. The results are presented in Fig. 16(b). Both solvers give substantial variations across the spectral range. Tidy3D shows internal consistency with variation below 3 between the 20 nm and 50 nm spectra. In Lumerical FDTD, the internal deviation is up to 9. The strong variations across the spectral range can also be seen in the mode crosstalk spectra for both solvers presented in Fig. 16(d). Lumerical FDTD shows an internal deviation of around 1 dB in the mode crosstalk, indicating that special attention should be taken when interpreting results from different bandwidth settings.
When using the custom non-uniform mesh in Lumerical FDTD, the mesh size is determined by the lower bound of the wavelength range, the highest refractive index in the simulation domain, and the specified number of cells per wavelength. It causes a small difference in the mesh size: 29.4 nm for 20 nm source bandwidth and 29.2 nm for 50 nm source bandwidth. By changing the mesh size to a constant value of 29 nm, the discrepancy between simulations with different bandwidths can be significantly reduced to below 5, as shown in Fig. 17(a) and (b). Comparing to the transmission spectra from Tidy3D in Fig. 16(b), there is still a large discrepancy over 16 due to mismatched positions of peaks and valleys.
3.5 Polarization splitter rotator
The PRS is identical to the polarizationsplitterrotator component in GDSFactory generic PDK. The device converts the injected fundamental TM mode into higher order TE mode and efficiently couple it to the fundamental TE mode of another waveguide. It can be used for polarization (de-)multiplexing in telecommunications. In this design, the component is defined in 220 nm thick Si, with SiO2 as bottom cladding and Si3N4 as top cladding. The refractive index of Si3N4 is approximated to 2.0 in the simulation bandwidth, as presented in [16]. Fig. 18 sketches the PSR with its key dimensions and Fig. 19 shows the component imported into Lumerical FDTD and Tidy3D. The input port on the left and the two output ports on the right are extended with 10 m long, 500 nm wide waveguides before importing into Lumerical FDTD and Tidy3D. A mode source is added on the left that inject fundamental TM mode at 1550 nm into the PSR. Fig. 20 presents the electric field intensity computed by the two solvers at a high resolution of 25 cells per wavelength, both showing efficient coupling from the left port to the upper right port. Tidy3D shows some standing wave patterns over the entire component.
The PSR is simulated at a range of resolutions from 6 to 25 cells per wavelength. Table 10 lists the runtime of the two solvers and the peak memory usage of Lumerical FDTD. Again, Tidy3D completes the simulations faster than Lumerical FDTD. In this case, the simulation domain size is 73.28.94 m3. For the highest resolution of 25 cells per wavelength, Tidy3D takes 21 minutes, while Lumerical FDTD needs 6 hours with 24 GB of memory usage. The total number of grid points and the elapsed simulation time and iterations are listed in Table 11. Tidy3D shows finer discretization and elapses almost two times longer simulation time.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 22 | 303 | 0.66 |
10 | 66 | 1073 | 2.14 |
15 | 285 | 3478 | 6.05 |
20 | 604 | 9665 | 13.06 |
25 | 1262 | 21896 | 24.01 |
Tidy3D | Lumerical FDTD | |
Total number of grids | 4.38108 | 2.63108 |
Number of grids in x | 4182 | 4152 |
Number of grids in y | 524 | 463 |
Number of grids in z | 200 | 137 |
Elapsed simulation time (s) | 2.5010-12 | 1.2210-12 |
Elapsed iterations | 74879 | 36980 |
Fig. 21(a) plot the transmission to the fundamental TE mode at the upper right port at 1550 nm as a function of resolution, simulated with a source bandwidth of 20 nm. Tidy3D shows a gradual increase from 5.1 at 6 cells per wavelength to 90.1 at 15 cells per wavelength. In contrast, Lumerical FDTD starts from 14.7 at 6 cells per wavelength, followed by a dip to 0.0, and then sharply increases to 94 at 15 cells per wavelength. Both solvers converge to a stable transmission around 95 but at different resolutions: Lumerical FDTD at 15 cells per wavelength and Tidy3D at 20 cells per wavelength. The difference between their stable value is below 0.5. Fig. 21(c) shows the mode crosstalk (transmission to the fundamental TM mode at the upper right port) at 1550 nm. The two solvers only agree above 20 cells per wavelength. These results suggest that the two solvers agree well when the resolution is sufficiently high, and Lumerical FDTD converges slightly faster in the simulation of the PSR.
Next, the simulation is examined with different source bandwidths at a resolution of 20 cells per wavelength, where the transmission in both solvers converges. Fig. 21(b) presents the transmission spectra computed by Tidy3D and Lumerical FDTD. Tidy3D shows excellent internal consistency with less than 0.1 variation between the results with 20 nm and 50 nm bandwidths. However, Lumerical FDTD gives significantly different spectra between these two bandwidth settings. The spectrum with 20 nm bandwidth peaks around 1545 nm with a maximum of 93.9, similar to Tidy3D results with an offset below 0.7. Meanwhile, the one with 50 nm bandwidth reaches a maximum of 96.8 at 1565 nm. The peak wavelength is shifted by 20 nm. In the mode crosstalk spectra presented in Fig. 21(d), there is also a considerable internal deviation for Lumerical FDTD. A minimum of -56.9 dB can be found around 1.56 m in the spectrum with 50 nm bandwidth. As the material dispersion and frequency dependent mode behavior are properly modeled in both solvers, it suggests an underlying problem with broadband simulations in Lumerical FDTD.
Similar to the analysis of the mode converter, there is a small difference of mesh size caused by varied source bandwidths in Lumerical FDTD. Given the resolution of 20 cells per wavelength, the mesh size is 22.1 nm and 21.9 nm for 20 nm and 50 nm bandwidths, respectively. By changing the mesh size to a constant value of 22 nm, the discrepancy between simulations with different bandwidths is almost eliminated, as shown in Fig. 22(a) and (b). The difference in transmission is reduced to 0.06. For the mode crosstalk in Fig. 22(b), the difference is 0.5 dB and the minimum moves to 1.54 m and increases to -48.1 dB. Comparing the transmission spectra in Fig. 22(a) to spectra from Tidy3D in Fig. 21(b), the inter-solver discrepancy is reduced to lower than 2 over the 50 nm bandwidth.
3.6 Ring resonator
The ring resonator uses the default layout of ringsingle in GDSFactory generic PDK. This component can serve as an optical filter. As sketched in Fig. 23, it consists of a 4.0 m-long directional coupler at the bottom with a gap width of 0.2 m, two 0.6 m-long vertical straight waveguides on the left and right, a 4.0 m-long horizontal straight waveguide at the top, and four bends connecting them with a radius of 10 m. The waveguide width is 500 nm. Fig. 24 shows the component imported into Lumerical FDTD and Tidy3D, with the input and output ports extended with 10 m long, 500 nm wide waveguides. A mode source is added on the left that inject the fundamental TE mode into the ring resonator with source bandwidths of 20 nm and 50 nm centered at 1550 nm. Fig. 25 shows the electric field intensity computed by the two solvers at 1550 nm for a resolution of 25 cells per wavelength, where the intensity distribution differs a little between the two solvers.
We simulate the ring resonator for spatial resolutions varied from 6 to 25 cells per wavelength. The simulation domain size for the ring resonator is 3224.054 m3. The runtime in the two solvers and the peak memory usage in Lumerical FDTD are listed in Table 12. As for previous components, Tidy3D is faster than Lumerical FDTD. For instance, Tidy3D completes the highest resolution simulation in 42 minutes, while Lumeical FDTD requires 28 hours on our workstation. Table 13 lists the number of grids and the elapsed simulation time and iterations at the highest resolution of 25 cells per wavelength. Tidy3D applies slightly finer discretization along all three dimensions.
Spatial resolution | Runtime, | Runtime, | Memory usage, |
(cells per wavelength) | Tidy3D (s) | Lumerical FDTD (s) | Lumerical FDTD (GB) |
6 | 135 | 1074 | 3.14 |
10 | 236 | 4527 | 8.36 |
15 | 971 | 16524 | 18.50 |
20 | 2414 | 44034 | 32.62 |
25 | 2503 | 100421 | 49.42 |
Tidy3D | Lumerical FDTD | |
Total number of grids | 4.68108 | 2.90108 |
Number of grids in x | 1873 | 1826 |
Number of grids in y | 1358 | 1302 |
Number of grids in z | 184 | 122 |
Elapsed simulation time (s) | 6.4010-12 | 6.4010-12 |
Elapsed iterations | 1.92105 | 1.95105 |
The simulation results are presented in Fig. 26. The monitors in these simulation has a wavelength step size of 0.2 nm. We then interpolate the transmission spectra from the output port with a step size of 0.02 nm for more accurate estimation of the lowest resonance wavelength in the simulation bandwidth and the corresponding full width at half maximum (FWHM, ). Also, we calculate the quality factor, namely the Q-factor, as following:
As presented in Fig. 26(a), Lumerical FDTD gives a resonance wavelength around 1545.6 nm for resolutions higher than 10 cells per wavelength, except for 15 where an abrupt increase occurs. Meanwhile, Tidy3D shows a gradual decrease in resonance wavelength and converges to around 1545.0 nm at 20 cells per wavelength. The corresponding FWHM is included in Fig. 26(b). Lumerical FDTD shows an increase from 0.84 nm at 6 cells per wavelength to 0.88 nm at 25 cells per wavelength with fluctuations in between, while Tidy3D maintains in the range between 0.88 nm and 0.9 nm. The difference between the two solvers reaches a stable value of 2 above 20 cells per wavelength. Similar trends can be found for the Q-factor presented in Fig. 26(c), where Lumerical FDTD generally decreases from 1839.4 to 1754.7 and Tidy3D oscillates between 1756.5 and 1714.9. The inter-solver difference in Q-factor also reaches 2 for resolutions higher than 20 cells per wavelength.
We then evaluate the performance over different source bandwidths. The transmission spectra from the two solvers for 20 nm and 50 nm bandwidth are shown in Fig. 26(d). When the source bandwidth is increased from 20 nm to 50 nm in Lumerical FDTD, the resonance wavelengths shift by 3.5 nm in Lumerical FDTD, which is about 47 of the free spectral range (FSR) of around 7.4 nm. In contrast, the resonance wavelengths shift up by 0.4 nm in Tidy3D, around 5 of the FSR of 7.6 nm. Similar to the discussions in previous sections, it suggests a potential impact of mesh size variation induced by custom non-uniform mesh and changed simulation bandwidth. This is proven by switching to a uniform mesh size of 22 nm over all three dimensions, where the difference is significantly reduced as shown in Fig. 27.
4 Discussion
In the previous section, we present the simulation results of a selection of passive components on SOI platform sourced from GDSFactory generic PDK, including directional coupler, waveguide crossing, 22 MMI, mode converter, a polarization splitter rotator (PSR), and a ring resonator. Through the simulation of directional coupler and waveguide crossing, we evaluate the propagation and coupling of fundamental modes. The multimode behavior is examined by the simulation of MMI. Moreover, the simulation of mode converter and PRS demonstrates higher-order mode coupling and polarization transitioning. Lastly, we compare the resonance behavior of ring resonator between the two solvers. The minimum feature size across these components is 200 nm, occurring in the MMI between the tapers and in the ring resonator between the ring and the straight waveguide. In general, Tidy3D shows finer discretization given the same spatial resolution setting and elapses longer simulation time before reaching the auto-shutoff threshold.
The key results are summarized in Table 14. A spatial resolution of 15 cells per wavelength is sufficient for most components in the two solvers. It corresponds to a mesh size of approximately 30 nm, as defined by the wavelength (1550 nm) and the refractive index of Si. The simulation of ring resonator is more demanding and requires a higher resolution of 20 cells per wavelength. On one hand, in the simulation of waveguide crossing, Tidy3D reaches a stable output at a lower resolution (10 cells per wavelength) than Lumerical FDTD. On the other hand, in the simulation of mode converter and PSR, Lumerical FDTD achieves stable results at lower resolutions than Tidy3D.
Component | Metrics | Lumerical FDTD | Tidy3D | ||
diectional coupler |
15 | 15 | ||
Simulation time | 22 min | 66 sec | |||
Bandwidth-induced variation | 3 | 0.2 | |||
Inter-solver discrepancy | 3.7 | ||||
waveguide crossing |
15 | 10 | ||
Simulation time | 4 min | 18 sec | |||
Bandwidth-induced variation | 0.1 | 0.03 | |||
Inter-solver discrepancy | 0.2 | ||||
22 MMI |
15 | 15 | ||
Simulation time | 15 min | 45 sec | |||
Bandwidth-induced variation | 0.05 | 0.05 | |||
Inter-solver discrepancy | 0.5 | ||||
mode converter |
10 | 15 | ||
Simulation time | 60 min | 7 min | |||
Bandwidth-induced variation | 9 | 3 | |||
Inter-solver discrepancy | 16 | ||||
PRS |
15 | 20 | ||
Simulation time | 58 min | 10 min | |||
Bandwidth-induced variation | 6 | 0.1 | |||
Inter-solver discrepancy | 2 | ||||
ring resonator |
20 | 20 | ||
Simulation time | 734 min | 40 min | |||
Bandwidth-induced variation | 47 | 0.5 | |||
Inter-solver discrepancy | 2 |
The two solvers generally agree well in a moderate simulation bandwidth of 20 nm over all examined components. The inter-solver discrepancy is quantified by examining the transmission to the target mode of each component, except for the ring resonator where we take the difference in FWHM and Q-factor. The difference is negligible for waveguide crossing and 22 MMI, and below 4 for directional coupler, PRS, and ring resonator. A higher difference of 16 is observed in the simulation of mode converter, a more demanding task where both solvers show considerable fluctuations at mismatched positions.
A change in the source bandwidth affects the two solvers differently. Tidy3D shows little variation as the bandwidth changes. However, when the bandwidth is increased from 20 nm to 50 nm, Lumerical FDTD sometimes gives different results, especially for the more complex components like the mode converter, PRS, and ring resonator. This variation can be reduced significantly by switching from custom non-uniform meshing to a uniform mesh. The former depends on not only the specified number of cells per wavelength but also the shortest wavelength in the simulation bandwidth. Note that the bandwidth-induced variation for ring resonator quantifies the shift of resonance wavelength as a percentage of FSR, while other components are evaluated based on the transmission to the target mode.
It worth noting that Tidy3D encountered the problem of diverging simulation in the simulations of directional coupler at higher resolutions (20 and 25 cells per wavelength) and mode converter across all resolutions. This is overcome by changing the boundary condition from PML to “absorber”. Boundary conditions do not need to be changed for our simulations in Lumerical FDTD.
The Tidy3D simulations presented in the previous sections are executed using version 2.7.7. During the course of our work, a newer version 2.8 has been released. It was brought to our attention that a new scheme named subpixel averaging is added so that the solver can better account for geometric features smaller than the mesh [23], functionally similar to CMT in Lumerical FDTD mentioned in Section 2. We examined the simulations of the directional coupler again using version 2.8.5 and subpixel averaging. As shown in Fig. 28, transmission to the target mode at 1550 nm converges to around 47 at a lower spatial resolution of 10 cells per wavelength.
Both Lumerical FDTD and Tidy3D are proven, reliable solvers for FDTD simulations. Lumerical FDTD is a mature solver offering extensive features and configurable settings. It can run efficiently on local hardware. The software includes a comprehensive example library and a scripting language. It also offers Python plugins (e.g. lumapi, lumopt). It can also be used in combination with other multiphysics solvers and circuit simulator offered by Lumerical. Tidy3D uses cloud-native GPU acceleration for rapid simulations, features a clear Python API and extensive examples, and is under active development to enhance its multiphysics and circuit simulation capabilities.
5 Conclusions
We presented a systematic comparison between Lumerical FDTD and Tidy3D across a range of passive components on SOI platforms representing different simulation scenarios. Our study focused on quantitative metrics including simulation runtime, memory consumption, bandwidth sensitivity, and inter-solver discrepancy. Under matched simulation conditions, the two solvers demonstrated excellent agreement in simulation results. Additionally, for most components, both solvers showed minimal variation in the output when the simulation bandwidth is increased from 20 nm to 50 nm. These results reinforce the reliability of both solvers.
Funding This work was supported by Natural Sciences and Engineering Research Council of Canada (NSERC, grant number RGPIN-2024-05029) and Mitacs through the Mitacs Elevate program.
Acknowledgment The authors thank Wesley Sacher for insightful discussions about the selection of devices and materials in this work. The authors also thank Flexcompute for providing information on the latest update and computation hardware of Tidy3D.
Disclosures The authors declare no conflicts of interest.
Data Availability The Python code underlying the results presented in this paper are available in Ref. [14].